# Learning to utilize image second-order derivative information for crisp edge detection

### v3

@@ IV-D Comparison to some other SOTA methods
As shown in Fig. 10, (d) and (e) are the output edge maps from DRC-RGB and Ours-RGB, respectively. The edge maps generated from our method exhibit consistent performance with those of predictions on BSDS500. The predictions from our method are cleaner and crisper than the top edge detector DRC, which demonstrates our effectiveness.
Table V illustrates that our method achieves impressive performance trained only on RGB images and HHA images, respectively. When we average the RGB version and HHA version, there is a noticeable improvement in S-Eval. The ODS achieves 0.768 and OIS achieves 0.780, which reaches the SOTA level. In crispness evaluation, our method obtains the best performance among other advanced methods. More precisely, ours-RGB is 32.8% and 22.9% higher than DRC-RGB in ODS and OIS, 31.9% and 27.3% higher than BDCN-RGB in ODS and OIS. Such a substantial increment fully demonstrates the effectiveness of our method in terms of edge crispness on the NYUD-V2 dataset. The Precision-Recall curves in Fig. 9 show that our method using EfficientNetV2 as the encoder has a top performance. These comparison results demonstrate the effectiveness of our method adequately.
TABLE VI: Quantitative comparison results on BIPED dataset. MS refers to multi-scale testing. Ours refers to our method based on EfficientNetV2.
Methods ODS OIS
RCF 0.849 0.861
BDCN 0.890 0.899
CATS 0.887 0.892
DexiNed 0.895 0.900
Ours 0.902 0.908
Ours-MS 0.903 0.909
@@
BIPED: Finally, we further evaluate our method on the BIPED dataset. Since the resolution of each image is relatively high ($1280\times 720$), we crop each image-label pair randomly into $320\times 320$ for training. There are currently limited methods that utilize the BIPED dataset due to its recent proposal. Therefore, we adopt four methods using deep learning for comparison which consists of RCF [10], BDCN [14], CATS [39], and DexiNed [30]. The quantitative comparison results are listed in Table VI and these results are similar to the performance shown on the BSDS500 and NYUD-V2. Our method obtains the highest performance on this benchmark. This is the first method to exceed 0.9 (ODS=0.903, OIS=0.909), demonstrating the powerful robustness of our method.
V Conclusion
@@
In this work, we propose a simple yet effective U-shape network named LUS-Net for crisp edge detection. We leverage the second-order derivative information to help the model locate true edge pixels more accurately. In addition, we construct a novel hybrid focal loss function to solve the issue of imbalanced pixel distribution. We address the issue of edge thickness and our method can generate crisp and clean contours without any post-processing. The experiment results show that we achieve state-of-the-art performance on three standard benchmarks, which demonstrate the advantages and effectiveness of our method. However, there is still room for improvement in our method. We still rely on pre-trained backbones on ImageNet which results in a cumbersome training process. Therefore, we will focus on exploring how to train the network from scratch in the future.

### v4

@@ 4.4 Comparison with SOTA methods
As shown in Fig. 10, (d), (e), and (f) are the output edge maps from EDTER, RankED, and our LUS-Net, respectively. The edge maps generated from our method exhibit consistent visualization with those of predictions on BSDS500. The predictions from our method are cleaner and crisper than the top edge detector EDTER and RankED, which demonstrates the effectiveness of our method.
Table 4 illustrates that our LUS-Net achieves strong performance across different image modalities (RGB, HHA, and RGB-HHA) and evaluation protocols (S-Eval and C-Eval). For RGB input under S-Eval, although our method achieves slightly lower scores (ODS=0.757) compared to recent methods like RankED (ODS=0.780) and EdgeSAM (ODS=0.783), it maintains competitive performance while using significantly fewer parameters (70.41M vs. 114.4M for RankED). More importantly, our method shows remarkable robustness in the challenging C-Eval protocol. For RGB input, LUS-Net achieves the ODS score of 0.540, significantly outperforming all recent SOTA methods including EDTER (0.416), RankED (0.477), and GCB-Net (0.456). This substantial performance gap (29.8% higher than EDTER, 13.2% higher than RankED, and 18.4% higher than GCB-Net) demonstrates our method’s enhanced capability in generating crisp edge maps. Similar robustness advantages can also be observed in HHA (0.435) and RGB-HHA (0.513) modalities, where our method consistently maintains superior performance across different image types. Specifically, LUS-Net achieves ODS=0.717 in HHA and ODS=0.768 in RGB-HHA under S-Eval protocol, while maintaining strong robustness in C-Eval with ODS=0.435 in HHA and ODS=0.513 in RGB-HHA modalities. In terms of computational efficiency, LUS-Net achieves 34.26 FPS on an A40 GPU, which is significantly faster than transformer-based methods including EDTER (2.1 FPS) and Boosting-ED (2.1 FPS). The Precision-Recall curves in Fig. 9 show that our LUS-Net has a top performance. These results demonstrate that our method achieves an excellent balance between accuracy, robustness, and efficiency.
@@
BIPED: Finally, we further evaluate our method on the BIPED dataset. Since the resolution of each image is relatively large ($1280\times 720$), we crop each image-label pair randomly into $320\times 320$ for training. We adopt recent SOTA methods for comparison which consists of SED [80], HED [24], RCF [25], CED-ADM [81], BDCN [29], PiDiNet [33], CATS [54], FCL-Net [55], DexiNed [45], EDTER [57], and Li et al. [82]. The quantitative comparison results are listed in Table 5 and these results are similar to the performance shown on the BSDS500 and NYUD-V2. Specifically, our LUS-Net achieves ODS=0.902, OIS=0.908, and AP=0.912, surpassing most existing methods. When employing multi-scale testing, the performance further improves to ODS=0.903, OIS=0.909, and AP=0.930. Notably, LUS-Net is the first method to exceed 0.9, demonstrating its powerful performance. Furthermore, compared to other SOTA methods including FCL-Net, EDTER, and DexiNed, our approach achieves more efficient architectural design while maintaining superior performance. These results further validate the effectiveness and generalization ability of our method across different datasets and scenarios.
Table 5: Quantitative comparison results on BIPED dataset. $\ddagger$ refers to multi-scale testing.
Methods ODS OIS AP Params
SED [80] 0.717 0.731 0.756 -
HED [24] 0.829 0.847 0.869 14.7M
RCF [25] 0.849 0.861 0.906 14.8M
CED-ADM [81] 0.810 0.835 0.869 -
BDCN [29] 0.890 0.899 0.934 16.3M
PiDiNet [33] 0.868 0.876 - 0.71M
CATS [54] 0.887 0.892 0.817 -
FCL-Net [55] 0.895 0.900 - 16.46M
DexiNed [45] 0.895 0.900 0.927 35.2M
EDTER [57] 0.893 0.898 - 468.84M
Li et al. [82] 0.890 0.894 0.902 21M
LUS-Net 0.902 0.908 0.912 70.41M
[…]
5 Conclusion
@@
Discussion: In this work, we proposed a simple yet effective U-shape network named LUS-Net for crisp edge detection. We leveraged the second-order derivative information to help the model locate true edge pixels more accurately. In addition, we constructed a novel hybrid focal loss function to solve the issue of imbalanced pixel distribution. Our method could generate crisp and clean contours without any post-processing, addressing the problem of edge thickness. The experiment results showed that we achieve state-of-the-art performance on three standard benchmarks, demonstrating the advantages and effectiveness of our method.
Limitation: There is still room for improvement. We currently rely on pre-trained backbones on ImageNet, which results in a cumbersome training process. Therefore, in future work, we will focus on exploring how to train the network from scratch, which could lead to further performance gains and a more efficient training pipeline.